OpenStudy (anonymous):

The coordinates of the vertices of quadrilateral ABCD are A(6, 7), B(4, 2), C(1, 5), and D(3, 8). ABCD is reflected across the x-axis to form PQRS. What is the length of diagonal PR?

OpenStudy (anonymous):

Since they are reflected, PR = AC. Find the distance between AC

OpenStudy (anonymous):

Using the distance formula? And that would be my answer?

OpenStudy (anonymous):

yes, find the distance from (6,7) to (1,5)

OpenStudy (anonymous):

i put the sqrt of -29 in my calculator and it came up as 'domain error' Help?

OpenStudy (anonymous):

how did you get sqrt -29? You can't take the root of a negative

OpenStudy (anonymous):

It's undefined. using pythagoras theorem: a^2 +b^2 = c^2, 2^2 + 5^2 = c^2 c^2 = 4 +25 = 29 c= sqrt 29

OpenStudy (anonymous):

i put in the equation. (1-6)^2 + (5-7)^2 then find the square root.

OpenStudy (anonymous):

after squaring, it should be positive then

OpenStudy (anonymous):

Oh i understand. Thank you!

OpenStudy (anonymous):

you're welcome